a.Caffeine has a therapeutic index of approximately 100:1. If the ED50 for caffeine is 12 mg/kg, what is the LD50?

b,Heroin has a half-life of approximate 10 minutes. Ronaldo injects 10mg of heroin IV. How much heroin will remain in his system 20 mins after injection?

c.Drug A has an ED50 of 10mg/kg and an LD50 of 100 mg/kg what is the therapeutic index of this drug? d

d.Drug X has an ED50 of 2 mg/kg and an LD50 of 400mg/kg. What is the therapeutic index for drug X?  

e. At the LD 50 of a drug what % of the subjects would die?
Answer #1

a. Therapeutic index=LD 50/ED 50

Here LD 50/ED 50=100/1

LD50/12=100/1

LD 50=1200 mg/kg

b.half life of heroin=10 min

Initial dose=10 mg IV

After 20 minutes remaining amount of heroin in the system=?

Since half life of the drug, is the time in which drug remains 50% of it's initial amount

So after 10 minutes the amount remain will be= 5 mg

In next 10 minutes the amount remain will be= 2.5 mg

Hence after total 20 minutes 2.5 mg of heroin will remain in the system.

c.Therapeutic index= LD50/ED50

=100/10= 10

d .TI=400/2= 200

e.At LD 50 of drug, 50% or half of the subjects will die